# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v3

@@ 3.4 Is the language model being truthful?
Table 1 illustrates the differences between the extracted in-context prior distribution (Section 2.3.2) and the LLM-elicited distribution (Section 2.2). No dataset has similar distributions according to their energy distance, demonstrating that the language model uses a different prior distribution to make its own predictions, to the one it provides for us to construct a separate predictive model.
These conclusions are supported in Appendix A.7, where we present histograms of the parameter values for each feature, which show that the two distributions visually differ significantly. This suggests that either the language model is unable to provide us with a reliable representation of its own prior knowledge or that it is unable to reliably use its own prior knowledge when making predictions. Since the in-context prior distributions are often centred around 00, we suspect that it is the latter and that the language model does not effectively use its own prior beliefs about a task in-context.
Since these priors disagree, we will investigate them further in Section 3.5, where we explore empirically whether in-context learning is approximating Bayesian inference and in Section 3.6, where we study decision criteria to understand which of these distributions is more likely to have generated the observed data using the Bayes factor.
3.5 Is in-context learning approximating Bayesian inference?
It is important that the model performs Bayesian inference, so that newly observed data points and prior knowledge are combined to form a posterior predictive distribution. In this section, we empirically evaluate the extent to which the language model is performing Bayesian inference in-context. This is done by measuring the difference between the extracted in-context posterior distribution and Monte Carlo (MC) posterior samples of the extracted prior in-context distribution (Section 2.3.4).
@@ 3.5 Is in-context learning approximating Bayesian inference?
Figure 6: The different posterior samples Here, we compare the distribution of the posterior extracted from the language model’s in-context predictions against the distribution that we sample using Monte Carlo methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined. For the diabetes dataset we have removed the long tails for improved visualisation.
The results of this analysis are given in Figure 6, which shows the extent to which the posterior distribution extracted from the language model’s in-context model agrees with Monte Carlo sampling on the extracted in-context prior. In this figure, we present the results for those datasets which achieve a low approximation error when estimating the posterior in-context parameters (Figure 5). The results of the other datasets are given in Appendix A.8, alongside those on other dataset splits.
The results for the two datasets given in Figure 6 present contradictory conclusions. For the Diabetes task, the posterior parameter samples from the in-context posterior and MC sampling on the in-context prior visually agree and produce a low energy statistic, suggesting that the LLM was able to approximate Bayesian inference in-context. On the other hand, for the Heart Disease task, we find that the two methods of sampling posterior parameters yielded different results, implying that for this task, the LLM was not correctly approximating Bayesian inference.
These results provide empirical evidence that whether in-context learning approximates Bayesian inference depends on the task. The LLM inconsistently applies the Bayes rule in context, which suggests that it should not be considered a reliable classifier. In this case, our prior elicitation framework presents itself as the more dependable option.
As the field develops, it is feasible that future LLMs can reliably perform Bayesian inference, and so in these cases we require a model selection tool to decide between different methods for incorporating a language model’s knowledge into a predictive task.
3.6 Bayes factor and model selection
@@
Figure 7: The distribution of the prior marginal likelihood, and the Bayes factor. This shows the distribution of prior likelihoods for the tasks. These were calculated using $25$ data points from each dataset, over $5$ test splits and $100$ samples of prompts. The Bayes factor is then calculated between prior elicitation and in-context learning, with a number larger than $1$ favouring prior elicitation. For the regression tasks, we use the log-scale and exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles.
To compare prior elicitation with in-context learning, when in-context learning is approximating Bayesian inference, we can calculate the Bayes factor as described in Section 2.4.
In the case of prior elicitation, we test the model as described in Section 2.2 and Equation 1. We approximate the marginal log-likelihood of this model by sampling the prior predictive log-likelihood of $\mathcal{D}$. For this, we use the same prior distribution $p_{M,T}(\theta)$ presented in 3.1 and sample $500$ predictions for $25$ data points, repeated with $5$ different splits of the dataset, to get samples of $p(\mathcal{D}|\alpha_{0})$
For in-context learning, we do not have access to $\varphi$ of the internal predictive model. However, since the language model provides us with predictions on targets given features directly, we can still approximate $p(\mathcal{D}|\alpha_{1})$ by repeated sampling of predictions. Here, we prompt the language model with our $100$ task descriptions and ask it to predict on $25$ data points from our dataset. We repeat this with $5$ different splits of the dataset to calculate the mean and standard deviation of the Bayes factor.
Figure 7 shows the distribution of the marginal likelihood for in-context learning, prior elicitation, and an uninformative prior along with the calculated Bayes factor (where reliably defined).
In the case of predicting Breast Cancer and Hypothyroidism, prior elicitation achieves a greater mean likelihood than both an uninformative prior and in-context learning. For these tasks, the Bayes factor and distribution of likelihoods suggest the use of prior elicitation, since it better describes the data (in terms of likelihood) and is cheaper to use for inference.
The Heart Disease prediction task favoured in-context learning, which provided a higher mean likelihood. For this task, the LLM returned an uninformative prior for the elicited distribution (Section 3.1), and used a prior with a smaller parameter range in-context (Section 3.3). In this case, the final choice of approach should depend on the setting, as using in-context learning for predictions might be too costly or unrealistic. Unlike the Bayes factor, the Bayesian Information Criterion (BIC) technique accounts for the computational complexity of classifiers (Section 2.4), but due to the large parameter sizes of an LLM and the availability of APIs, it could unfairly disfavour in-context learning. This approach may be more informative for some settings, and for Heart Disease prediction, would recommend prior elicitation.
For regression tasks, prior elicitation is preferred for $y=2x_{1}-x_{2}+x_{3}$, whilst for Diabetes, in-context learning produces greater log-likelihoods. However, here, both in-context learning and prior elicitation provide a high prior mean squared error and so a low marginal likelihood, making the Bayes factor unreliable to compute. Again, in the case of Diabetes prediction, the final model decision relies on the available resources.
Whether prior elicitation or in-context learning provides prior distributions with greater likelihoods depends on the task, with prior elicitation being the better choice over a uninformative prior for all datasets tested. However, when factoring in the methods’ abilities to estimate Bayesian inference and the cost of producing predictions, prior elicitation is likely the better option for the cases tested.
3.7 Choosing between priors
As seen in Appendices A.13.1 and A.13.2, the prior distributions elicited from language models do not always provide the best posterior performance. We consistently find that GPT-3.5-turbo and GPT-4-turbo provide better prior distributions than an uninformative prior; however, this may not be the case for all datasets or language models. Therefore, given prior distributions elicited from different language models, we would like to choose the best one for our predictive task. For this, given some collected data, we can use the Bayes factor (Section 2.4) as a model selection technique to decide between the prior distributions elicited from different language models. Alternatively, we could use other model selection techniques, such as the Bayesian Information Criterion. This is because we are performing the Bayesian inference ourselves, given the prior distributions. Furthermore, this method also allows us to compare the elicited distributions to the uninformative prior to ensure that the priors from our language models do not harm predictive performance. By using a temperature parameter, it is also possible to control the strength of the elicited prior, which can indicate our trust in the distribution.
@@ 4 Limitations and future research
In this work, we consider the language model as a domain expert to elicit prior distributions. These prior distributions need to be better than an uninformative prior to improve model performance and reduce data collection costs. A natural next step therefore is to question how these prior distributions compare to those elicited from domain experts, and whether language models provide an additional opportunity to reduce costs by supplementing expert elicited priors or by improving the expert elicitation framework. We began to explore how experts could work alongside an LLM by producing more detailed task descriptions (Section 3.2 and Appendix A.10), but other ways of integrating experts would be interesting to explore.
As priors elicited from language models are a product of their training, they could contain unwanted bias. An approach to avoid this would be to assign all features that could be considered sensitive to unwanted bias with an uninformative prior. However, prior distributions on other features might indirectly cause unwanted bias in a predictive model, and it is important to carefully think about whether these features should also be assigned an uninformative prior. Given an elicited prior distribution and observed examples, the Bayes factor can again be used for model selection to choose between combinations of an elicited prior with some or all features replaced with an uninformative prior. Additionally, by providing information in the task description, an expert could steer the language model away from unwanted bias in its provided distributions.
Finally, an interesting question is how we extend our results to neural networks. An immediate option is to ask a language model to suggest appropriate data augmentation techniques for a task, effectively reducing the cost of data collection by extending the current dataset. Another possible solution is to use the output of in-context predictions as additional features to be learnt from. Further exploration of methods in this direction would be beneficial.
5 Further results
In the Appendix, we supplement our findings with further results and discussions. Firstly, we present the complete prior elicitation and model selection process on a new dataset in Appendix A.5.5. Then we present additional results of prior elicitation and in-context learning (Appendix A.6, A.7, A.8), where we support the conclusions drawn in Sections 3. Next, since some of the datasets tested are in the public domain and might have appeared in the language model’s training, in Appendix A.9, we provide a discussion on whether the language model has memorised the datasets tested.
To further understand the language model’s behaviour to prompts, we study whether the language model is respecting our in-context model class specifications (Appendix A.3) and test how the detail of the task description affects the posterior performance of the elicited priors (Appendix A.10), as well as how experts can supplement the process in a controlled synthetic experiment. Further to this, we explore how the number of task descriptions affects the priors elicited (Appendix A.11).
Finally, to contextualise the value of prior elicitation from language models, we report the cost of performing prior elicitation (Appendix A.12) and evaluate the quality of priors elicited from other language models with different pricing tiers (Appendix A.13.1) as well as open-source models (Appendix A.13.2).
@@ 6 Conclusion
In this work, we presented a method for extracting task information from an LLM to build probabilistic linear models with fewer data points. By asking a language model to play the role of a domain expert, we were able to elicit prior distributions for predictive tasks, potentially reducing the cost of data collection and improving the accuracy of our models when we have access to few labelled examples (Section 3.1). As demonstrated in Appendix A.6.2, these prior distributions can significantly impact the delivery of care in healthcare settings, where label collection can be costly and time-consuming, but the need for accurate predictive models is great.
We find that the prior distributions elicited from GPT-3.5-turbo achieve better accuracy or mean squared error with fewer data points than an uninformative prior for a variety of tasks (Section 3.1). This performance improvement is most visible when few training examples were available, suggesting its use in environments where label collection is challenging. We also find that incorporating expert information into the task descriptions can enable further improvements in predictive accuracy, and that the language model updates its prior distributions as we would expect (Section 3.2 and Appendix A.10).
Since our approach to prior elicitation can also be seen as a direct replacement for in-context learning for datasets with numerical features, we present a model selection framework, based on the Bayes factor (Section 3.6). However, the Bayes factor assumes that the in-context learner is performing Bayesian inference. Therefore, we use maximum likelihood estimation and sampling distributions to approximate the internal model used during in-context learning and its prior and posterior distributions (Section 3.3). This allows us to inspect the truthfulness of the language model when we conduct prior elicitation (Section 3.4), as well as empirically question whether the language model is performing Bayesian inference on a given task (Section 3.5).
We find that prior elicitation is an improved alternative to in-context learning in our setting, as we can inspect priors, confidently perform Bayesian inference, and often achieve reduced error. Prior elicitation also extends the use of the language model’s prior knowledge of predictive tasks to previously infeasible situations, such as low-resource environments or where feature observations are frequent but labelled examples are few.
We have demonstrated that using language models for expert prior elicitation is a cost-effective method for improving the predictive performance of probabilistic linear models in environments with few labels and provides a competitive alternative to in-context learning.
@@
AC: Original idea conception, software, analysis of results, writing, editing, and reviewing, funding acquisition. RK: Original idea conception, supervision, analysis of results, reviewing. PB: Supervision, analysis of results, reviewing, funding acquisition.
@@ Acknowledgments
This study is funded by the UK Dementia Research Institute (UKDRI) Care Research and Technology Centre funded by the Medical Research Council (MRC), Alzheimer’s Research UK, Alzheimer’s Society (grant number: UKDRI–7002), and the UKRI Engineering and Physical Sciences Research Council (EPSRC) PROTECT Project (grant number: EP/W031892/1). Infrastructure support for this research was provided by the NIHR Imperial Biomedical Research Centre (BRC) and the UKRI Medical Research Council (MRC). Further funding was provided by the Turing scheme. The funders were not involved in the study design, data collection, data analysis, or manuscript writing.
@@ A.1 Availability of code
Within the supplementary materials, we provide code for generating all figures and results, as well as provide a pipeline for using our method on additional datasets.
Our code is implemented in python (v3.11.9) using pip (v24.0).
For our experiments, we use the Python packages:
  • blackjax (v1.2.3) [44]: For Monte Carlo sampling on the in-context prior samples.
  • jax (v0.4.31) [45]: For calculating some metrics and use with blackjax.
  • numpy (v1.26.4) [46]: For manipulation of arrays and computing mathematical equations.
@@
  • pandas (v2.2.2) [47]: For manipulating the datasets and results.
  • pymc (v5.17.0) [48]: For Monte Carlo sampling on the elicited priors.
  • scikit-learn (v1.5.1) [49]: For loading and preprocessing the data.
@@
A.2 Formal details on extracting the internal prior and posterior
@@
In this section, we provide a description of the method presented in Section 2.3.2 in more detail.
@@
Given a language model $M$ and task $T$ that provides in-context predictions, if we can construct a function that allows us to sample a single parameterisation $\varphi\sim p_{M,T}(\varphi)$, then we can estimate $p_{M,T}(\varphi)$ using a sampling distribution. Following Murphy [32], we define an estimator:
@@
A.3 Does the language model use the model class we specify?
Figure 8: In-context predictions for the task: $y=2x_{1}-x_{2}+x_{3}$ compared to the true values. Here, feature values are generated using $X\sim\mathcal{U}(-5,5)$ and the targets are calculated using $y=2x_{1}-x_{2}+x_{3}$, which is specified fully in the in-context prompt. The language model is then asked to make predictions on $X$. This shows the true targets $y$ against those predicted by the language model, as well as the binned root mean squared error.
As part of comparing prior elicitation with in-context learning, we use maximum likelihood sampling to extract the internal parameter distributions (discussed in Section 3.3), which we use to test whether the language model is: (1) providing a similar prior distribution through elicitation to the one it is using in-context, and (2) empirically test whether the language model is approximating Bayesian inference. However, when extracting the internal parameters of the in-context model, we assume that the language model is applying the model class that we specify in the task description.
However, we find that for some datasets, the language model does not use the same model in context as the one we specify. For the synthetic data in particular, we specify the relationship as $y=2x_{1}-x_{2}+x_{3}$. Here, we can measure whether the language model is using the specified model in-context to predict the target by calculating the error between the true $y$ values and those returned by the language model’s in-context predictions. This is done by measuring the mean squared error between the $y_{n}$ given in Equation 6 and the true $\hat{y}_{n}=2x_{n,1}-x_{n,2}+x_{n,3}$ calculated using the $x_{n}$ in Equation 6.

### v4

@@ 5.5 Is the language model consistent?
5.6 Does in-context learning use Bayesian inference?
It is important that in-context learning performs Bayesian inference, so that new data updates the LLM’s knowledge reliably. We empirically study this by comparing the approximated in-context posterior with the Monte Carlo (MC) posterior samples of the approximated in-context prior.
These two posterior samples are shown in Figure 6, where we focus on those datasets where in-context learning produced linear predictions, providing a low posterior approximation error. For the Diabetes task, samples from the approximated in-context posterior and MC sampling on the approximated in-context prior visually agree and produce a low energy statistic, suggesting the LLM is performing Bayesian inference in-context. However, for the Heart Disease task the two posterior samples differ, showing that the LLM inconsistently applies the Bayes rule in-context. Results of the other datasets and training splits are given in Appendix A.14. This suggests LLM-elicited priors are more reliable, particularly when access to the posterior distribution is important, such as for uncertainty estimation.
@@ 5.6 Does in-context learning use Bayesian inference?
Figure 6: The in-context posterior distribution. We compare the posterior extracted from GPT-3.5-turbo’s in-context predictions against the distribution that we sample using Monte Carlo methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined. For the diabetes dataset we have removed the long tails for improved visualisation.
It is feasible that future LLMs can reliably perform Bayesian inference in-context, requiring a model selection strategy to decide between in-context learning and LLM-elicited priors.
5.7 Using Bayes factor for model selection
If in-context learning approximates Bayesian inference, we calculate the Bayes factor between it and AutoElicit priors.
@@ 5.7 Using Bayes factor for model selection
Figure 7: The distribution of the prior marginal likelihood and the Bayes factor for GPT-3.5-turbo. This shows the distribution of prior likelihoods, calculated using $25$ data points from each dataset, over $5$ test splits and $100$ samples of prompts. The Bayes factor is calculated between prior elicitation and in-context learning, with a number greater than $1$ favouring prior elicitation. For the regression tasks, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles in the histograms.
Figure 7 shows the marginal likelihood distribution for in-context learning, AutoElicit, and an uninformative prior along with the Bayes factor (where reliably defined). Unlike in-context learning, AutoElicit priors are the better choice over an uninformative prior for all datasets tested. Whether AutoElicit or in-context learning provides better priors depends on the task. However, in only two of the five datasets tested, in-context learning reliably produced linear predictions, and in only one of those it approximated Bayesian inference. Considering this, the cost, and the ease of interpretability, AutoElicit is likely the better option.
This model selection strategy is designed to be more relevant as LLMs become more reliable, and their cost reduces.
@@ 6 Limitations and future research
This work explores how experts can guide AutoElicit priors with natural language, but studying other ways of incorporating experts into the framework would be interesting. For example, by fine-tuning LLMs on relevant literature.
As AutoElicit priors are a product of their training, they might contain unwanted bias. To avoid this, an approach would be to assign all features that are sensitive to unwanted bias, directly or indirectly, with an uninformative prior. Given an elicited prior and observed examples, the Bayes factor can then be used to choose between the elicited prior and an uninformative prior. Additionally, experts can steer LLMs away from unwanted bias through task descriptions.
Finally, it would be interesting to extend our results to neural networks. An immediate option is to build on the literature asking the LLM to suggest data augmentation techniques for a task, effectively reducing the cost of data collection by extending the current dataset. Further, extending our results to non-linear models such as random forests or Gaussian processes would be valuable.
@@ 7 Conclusion
This work shows that, by asking a language model to play the role of a domain expert, we are able to elicit prior distributions for predictive tasks, potentially reducing the cost of data collection and improving the accuracy of our models when we have access to few labelled examples.
We find the priors elicited from GPT-3.5-turbo, GPT-4-turbo, and DeepSeek-R1-32B-Q4 achieve better accuracy or mean squared error with fewer data points than an uninformative prior for a variety of tasks. This improvement is most visible when few training examples are available, suggesting its use in environments where label collection is challenging.
Adding expert guidance into the task descriptions enables further improvements to predictive accuracy, as the LLM updates elicited distributions as we would expect. AutoElicit can therefore also be used to simplify the typical expert prior elicitation process, by only requiring experts to provide insights as natural language. We find that prior elicitation with AutoElicit is an improved alternative to in-context learning in our setting, as we can inspect priors, confidently perform Bayesian inference, and often achieve reduced error.
On our private dataset we find a large and quantifiable performance improvement when predicting cases of UTI. Within our study on dementia care the resulting model created with LLM-elicited priors can provide better predictions of adverse health conditions, reliable uncertainty estimates, and remain interpretable and trusted by clinicians.
@@
AC: Original idea conception, software, analysis of results, writing, editing, reviewing, funding acquisition. RGK: Original idea conception, supervision, analysis of results, writing, editing, reviewing, funding acquisition. PB: Supervision, analysis of results, reviewing, funding acquisition.
@@ Acknowledgments
This study is funded by the UK Dementia Research Institute (UKDRI) Care Research and Technology Centre funded by the Medical Research Council (MRC), Alzheimer’s Research UK, Alzheimer’s Society (grant number: UKDRI–7002), and the UKRI Engineering and Physical Sciences Research Council (EPSRC) PROTECT Project (grant number: EP/W031892/1). Infrastructure support for this research was provided by the NIHR Imperial Biomedical Research Centre (BRC) and the UKRI Medical Research Council (MRC). Further funding was provided by the Turing scheme, UK, and resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. Additionally, RGK acknowledges support from a Canada CIFAR AI Chair. The funders were not involved in the study design, data collection, data analysis, or manuscript writing.
@@ A.1 Availability of code
Within the supplementary materials, we provide code for generating all figures and results, as well as a pipeline for using our method on additional datasets.
The open-source code makes it simple to perform expert prior elicitation using LLMs: {minted}[mathescape, numbersep=5pt, frame=lines, framesep=5mm, ]python # import the elicitation function from llm_elicited_priors.gpt import ( get_llm_elicitation_for_dataset )
# elicit priors for a dataset get_llm_elicitation_for_dataset( # the language model client # requires a .get_result(messages) client=llm_client, # the task descriptions # to iterate over system_roles=system_roles, user_roles=user_roles, # the dataset feature names feature_names=data.feature_names, ) We provide classes that wrap common LLMs (Appendix A.15), whilst making it easy to adapt to new models.
This code is implemented in python (v3.11.9), using pip (v24.0). For our experiments, we use the packages:
  • blackjax (v1.2.3) (Cabezas et al., 2024): For Monte Carlo sampling on the in-context prior samples.
  • jax (v0.4.31) (Bradbury et al., 2018): For calculating some metrics and use with blackjax.
  • numpy (v1.26.4) (Harris et al., 2020): For manipulation of arrays and computing mathematical equations.
  • ollama (v0.4.7): For loading the DeepSeek model.
@@
  • pandas (v2.2.2) (Wes McKinney, 2010): For manipulating the datasets and results.
  • pymc (v5.17.0) (Abril-Pla et al., 2023): For Monte Carlo sampling on the elicited priors.
  • scikit-learn (v1.5.1) (Pedregosa et al., 2011): For loading and preprocessing the data.
@@
A.2 Further details on the related work
A.2.1 Using LLMs to elicit expert priors
Choi et al. (2022) explores the use of an LLM to perform feature selection, causal discovery, and to modify the reward function in a reinforcement learning task. In the first case, they observe that when listing features from two datasets, GPT-3 can correctly identify the features that correspond to a single specified predictive task. Additionally, when applied to causal discovery, the language model is often able to predict the directionality of causal relationships in the Tüebingen Cause-Effect Pairs dataset (Mooij et al., 2016). This inspires us to explore the possibility of using LLMs for eliciting prior parameter distributions over linear models.
Gouk & Gao (2024) use LLMs to generate synthetic data to elicit priors on public datasets. A logistic regression model is then pre-trained using this, defining its prior. However, this work does not discuss the memorisation of public datasets by LLMs, which is shown to be substantial Bordt et al. (2024), supported by the results in Appendix A.8.
Further, Selby et al. (2024) elicit parameters of a beta distribution to define a prior for predictive tasks where results from human experts are available. The priors from four LLMs were visually compared to those elicited from human experts using the SHELF (Oakley & O’Hagan, 2019; Gosling, 2018) method, which demonstrated significant differences between experts and LLMs. However, they show, by considering prior predictive distributions, that GPT-4 (OpenAI et al., 2024) can provide parameter distributions that allow for zero-shot performance that compares to models trained on tens of data points. In our work, we extend this by using the LLM to elicit multiple Gaussian distributions as priors. We combine these to create a mixture model for a more general distribution over our parameters and test their utility for linear predictive models.
In addition, Requeima et al. (2024) prompt language models with task descriptions and collected data points to generate posterior predictive distributions for regression tasks, which they find outperform Gaussian processes in some tasks. This is done by repeatedly sampling a language model at a new feature value, given the previous input and output observations, and a given text description that contains prior information from a human expert.
Furthermore, Zhu & Griffiths (2024) describe the use of iterated learning to elicit prior distributions from language models for causal learning, proportion estimation, and predicting everyday quantities. The authors employ a Markov chain Monte Carlo method to access an LLM’s prior knowledge by performing successive inference using previous model outputs that tend towards a limiting distribution. In doing this, the authors assume that in-context learning is a form of Bayesian inference. However, in their work Zhu & Griffiths (2024) do not consider how prior elicitation from LLMs can be used for predictive models or test whether an LLM is performing Bayesian inference, both of which we study in this work.
A.2.2 In-context learning
Learning from demonstrations in-context has proven to be effective for several natural language tasks, demonstrating the surprising ability of language models to perform outside of their intended use (Brown et al., 2020; Dong et al., 2022; Hegselmann et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023). In Brown et al. (2020), the authors study GPT-3’s strong ability on natural language tasks after seeing a few examples, one example, or none at all. This work inspired others to study the extent to which language models can learn from demonstrations given as text alongside a task description. Further, Xie et al. (2021) study the mechanics of in-context learning and describes it from a Bayesian Inference perspective. They find that in-context learning succeeds when language models can infer a shared concept across individual examples in a prompt. Min et al. (2022) argue that in-context learning allows the language model to observe the label space, the distribution of input text, and the task format, as they find that swapping the real labels in demonstrations for random labels does not significantly reduce accuracy. Most of the existing work focuses on natural language tasks Dong et al. (2022), leaving numerical problems less explored and a focus of our work.
In Hollmann et al. (2022), the authors train a transformer model to solve synthetically generated classification tasks from a prior over tabular datasets. Given a new task and demonstrations, this transformer performs in-context learning to make predictions using inference steps only. Furthermore, Akyürek et al. (2022) and Von Oswald et al. (2023) discuss the types of learning algorithms that language models (and more generally, transformers) can implement in-context. Focusing on linear models, Akyürek et al. (2022) show that, by construction, transformers are capable of performing gradient descent and closed-form ridge regression. Akyürek et al. (2022) also demonstrate that trained in-context learners closely match predictors trained through gradient descent. Further, Von Oswald et al. (2023) show that transformers, by incorporating multilayer perceptrons into their architectures, can also solve non-linear regression tasks, and Dalal & Misra (2024) propose a Bayesian framework to reason about a language model’s internal state and suggest that in-context learning is consistent with Bayesian learning. In our work, we test empirically, through the approximation of the internal model used by a language model for its in-context predictions, whether for a given task, the language model is performing Bayesian inference before and after seeing demonstrations.
A.3 Formal details on extracting the internal prior and posterior
@@
In this section, we provide a description of the method presented in Section 4.2 in more detail.
@@
Given a language model $M$ and task $T$ that provides in-context predictions, if we can construct a function that allows us to sample a single parameterisation $\varphi\sim p_{M,T}(\varphi)$, then we can estimate $p_{M,T}(\varphi)$ using a sampling distribution. Following Murphy (2022), we define an estimator:
@@
A.4 Does the language model use the model class we specify?
As part of comparing AutoElicit with in-context learning, we use maximum likelihood sampling to extract the internal parameter distributions (discussed in Section 5.4), which we use to test whether the language model is: (1) providing a similar prior distribution through elicitation to the one it is using in-context, and (2) empirically test whether the language model is approximating Bayesian inference. However, when extracting the internal parameters of the in-context model, we assume that the language model is applying the model class that we specify in the task description.
Figure 8: GPT-3.5-turbo in-context predictions for the task: $y=2x_{1}-x_{2}+x_{3}$ compared to the true values. Here, feature values are generated using $X\sim\mathcal{U}(-5,5)$ and the targets are calculated using $y=2x_{1}-x_{2}+x_{3}$, which is specified fully in the in-context prompt. The language model is then asked to make predictions on $X$. This shows the true targets $y$ against those predicted by the language model, as well as the binned root mean squared error.
However, we find that for some datasets, the language model does not use the same model in context as the one we specify. For the synthetic data in particular, we specify the relationship as $y=2x_{1}-x_{2}+x_{3}$. Here, we can measure whether GPT-3.5-turbo is using the specified model in-context to predict the target by calculating the error between the true $y$ values and those returned by the language model’s in-context predictions. This is done by measuring the mean squared error between the $y_{n}$ given in Equation 7 and the true $\hat{y}_{n}=2x_{n,1}-x_{n,2}+x_{n,3}$ calculated using the $x_{n}$ in Equation 7.